2 diamond rings and 4 silver rings cost $1440. a diamond ring and a silver ring cost $660. how much does a silver ring cost ?
i know the answer is $60. we did it with process of elimination and could not figure out another way
we got 2x +4y=1440
x =y =660

Answer by mananth(16949)

let diamond ring be $x
Silver ring be $y
..
2x+4y=1440.................1
x+y=660.....................2
...
multiply (2) by -2 and add to (1)
-2x-2y=-1320
add to (1) we get
4y-2y=1440-1320
2y=120
y= $60 cost of silver ring.
....
m.ananth@hotmail.ca